# BAnG: Bidirectional Anchored Generation for Conditional RNA Design

### v1

@@ 5 Conclusion
This study introduces a novel deep-learning model, RNA-BAnG, and a sequence generation method, BAnG, for designing RNA sequences that bind to a given protein. Unlike previous approaches, our model demonstrates remarkable flexibility by eliminating the need for extensive structural or interaction data. Although our model relies on protein structure, AlphaFold predictions make sequence information sufficient. This innovation significantly broadens the applicability of our method, making it a versatile tool for RNA-protein interaction studies.
@@
The code and the model, along with the model weights, will be available upon publication.
@@ Acknowledgments
This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011015647 made by GENCI. A substantial part of the computations presented in this paper was performed using the GRICAD infrastructure (https://gricad.univ-grenoble-alpes.fr), which is supported by Grenoble research communities.
@@ Appendix A RNA-BAnG architecture
RNA-BAnG architecture is schematically presented in Figure 5 of the main text and is composed of protein and nucleotide modules, detailed in Figure A.1 and Figure A.2. The latent dimension of the model is $c_{s}=128$, all heads dimensions are set to $c_{h}=64$. Feedforward blocks have a scaling factor $n=2$. We set the number of protein and nucleotide modules to 10 each. While reducing this number led to poorer performance, increasing it did not yield any noticeable improvements. We adjusted the rest of the hyperparameters by selecting the smallest model size combination that ensured stable convergence. Resulting model contains 14,5 million parameters.
Figure A.1: Schematic illustration of the protein module in the RNA-BAnG architecture. Here, ${\bf r_{aa}}$ is the number of protein residues, $\bigotimes$ stands for concatenation.
Figure A.2: Schematic illustration of the nucleotide module in the RNA-BAnG architecture. Here, ${\bf r_{na}}$ is the number of nucleotides, ${\bf r_{aa}}$ is the number of protein residues, $\bigotimes$ stands for concatenation.
@@ B.1 Toy model
To avoid direct memorization on the synthetic task, we opted for a compact model, same for each tested method. Its final configuration is a two-block transformer (with a hidden dimension of 64) with the RoPE positional attention and two attention heads, resulting in 17k parameters. We trained the model for each method for 80k steps with a batch size of 8, using ADAM optimizer (Kingma & Ba, 2017) with the learning rate of 0.0001. During the inference, the length was fixed to 50 tokens for iterative methods.
@@ B.3 Generated synthetic sequences
Appendix C Data processing
@@ Appendix D Training details
As explained in the main text, we conducted the RNA-BAnG training in two steps. In the first step, we used a batch size of 64 and trained for 255k steps. In the second step, we used a batch size of 8 and trained for 216k steps. We stopped training when the validation loss decline became negligible. In both training steps, we used the ASMGrad (Reddi et al., 2019) variation of the ADAM optimizer (Kingma & Ba, 2017) with default parameters and the learning rate of 0.0001. Learning rate was warmed up linearly for the first 1k steps and then decayed exponentially with $\gamma=0.99$ and a period of 1k steps. Complete training took 4 days on a single MI120 AMD GPU.
Appendix E Evaluation details
@@ E.1 Test data
The RNA Compendium study provides 244 samples, each comprising a protein sequence paired with approximately 200,000 RNA sequences and their corresponding experimental binding scores. These samples are designated by the authors as RNCMPT00XXX, where XXX represents a numerical identifier. For simplicity, throughout this paper, we refer to these samples solely by their numerical identifiers XXX.
We processed the data as follows. For each sample, RNA sequences were ranked by their binding scores. The top 2,000 sequences were labeled as the positive (interacting) class, while the bottom 2,000 were labeled as the negative (non-interacting) class. To eliminate sequence redundancy and prevent data leakage, duplicate RNA sequences were removed using CD-HIT-EST with a sequence identity threshold of 90%. These sequences were then randomly split into training and testing sets, so that each test set would have a 1,000 sequences. We then trained individual DeepCLIP models for each sample, following the protocol outlined by the authors. Only models achieving an area under the receiver operating characteristic curve (AUROC) of 0.95 or higher on the corresponding test set were selected for further analysis.
@@ E.2 RNA-BAnG sampling
For RNA-BAnG, we sample tokens from their distributions using the top-k strategy with $k=4$. Maximum sequence length is set to 50 nucleotides. Average sampling time is 20 minutes for 1,000 RNA sequences. To sample random sequences, we first randomly select a sequence length between 40 and 50, then uniformly sample that many nucleotides. This length choice ensures that the mean sequence length of random samples matches that of RNA-BAnG-generated sequences (45.6 nucleotides).
Appendix F Additional results
@@
Figure F.1 shows the proportion of high-affinity generated sequences as a function of AlphaFold pLDDT scores and protein sequence similarity with the train set. Appendix G details protein sequence similarity calculations. Figure F.2 shows kernel density plots calculated over the whole test set of 71 samples. We assessed four different DeepCLIP score threshold values (0.65, 0.75, 0.85, 0.95, from left to right in the Figure). We can conclude that the value of 0.95 is too stringent, as the mean of the positive experimental sets approaches the value of 0.5. Subjectively, the most visually appealing threshold value is 0.75. Nonetheless, in the main text (Figure 6) we continuously assess all the values by plotting the mean proportions above a certain threshold value. Distributions of DeepCLIP scores for each of 71 test samples are depicted in Figure F.3 and Figure F.4.
@@
Figure F.3: Distribution of DeepCLIP scores for the first part of the test samples (RNAcompete sample IDs on top). Scores of RNA-BAnG are in green. Scores of positive and negative experimental sequences are in blue and red, respectively.
Figure F.4: Distribution of DeepCLIP scores for the second part of test samples (RNAcompete sample IDs on top). Scores of RNA-BAnG are in green. Scores of positive and negative experimental sequences are in blue and red, respectively.
Appendix G Tools parameters
@@
Clustering of nucleotide sequences was always performed using CD-HIT-EST (Li & Godzik, 2006) with a sequence identity threshold of 90%. We used CD-HIT-EST with these parameters: -c 0.9 -n 9 -d 0 -T 10 -U 10 -l 9. For protein sequence identity to the train set calculations we used blastp (Altschul et al., 1990) with default parameters. For the nucleotide sequences similarity to the train set calculations we used blastp (Altschul et al., 1990) with default parameters.

### v2

@@ 5 Conclusion
This study introduces a novel deep-learning model, RNA-BAnG, and a sequence generation method, BAnG, for designing RNA sequences that bind to a given protein. Unlike previous approaches, our model demonstrates remarkable flexibility by eliminating the need for extensive structural or interaction data during inference. Although our model relies on protein structure, AlphaFold predictions make sequence information sufficient. This innovation significantly broadens the applicability of our method, making it a versatile tool for RNA-protein interaction studies.
@@
The code and the model, along with the model weights, are available at https://github.com/rsklypa/RNA-BAnG. The code allows for running the RNA sequences generation process for any correct protein 3D structure as input.
Impact Statement
The main purpose of this work is to advance the field of generative models. However, the applications of this method may have social and industrial benefits. Potential applications include in-silico SELEX approaches, RNA vaccine design, the development of novel drugs, and some other pharmaceutical tasks.
@@ Acknowledgments
This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011015647 made by GENCI. A substantial part of the computations presented in this paper was performed using the GRICAD infrastructure (https://gricad.univ-grenoble-alpes.fr), which is supported by Grenoble research communities. The authors are very thankful to Yoann Fauconnet and Jessica Andreani from I2BC, CEA Paris-Saclay for their support and advise on RNA scoring models. The authors also acknowledge the support from the PHC MAIMONIDE 2023 program, project number 49285YJ.
@@ Appendix A RNA-BAnG Architecture
RNA-BAnG architecture is schematically presented in Figure 5 of the main text and is composed of protein and nucleotide modules, detailed in Figure A.1. The latent dimension of the model is $c_{s}=128$, all heads dimensions are set to $c_{h}=64$. Feedforward blocks have a scaling factor $n=2$. We set the number of protein and nucleotide modules to 10 each. While reducing this number led to poorer performance, increasing it did not yield any noticeable improvements. We adjusted the rest of the hyperparameters by selecting the smallest model size combination that ensured stable convergence. Resulting model contains 14,5 million parameters.
Figure A.1: Schematic illustration of the protein (on the left) and nucleotide (on the right) modules in the RNA-BAnG architecture. Here, ${\bf r_{na}}$ is the number of nucleotides, ${\bf r_{aa}}$ is the number of protein residues, $\bigotimes$ stands for concatenation.
@@ B.1 Toy Model
To avoid direct memorization on the synthetic task, we opted for a compact model, same for each tested method. Its final configuration is a two-block transformer (with a hidden dimension of 64) with the RoPE positional attention and two attention heads, resulting in 17k parameters. We trained the model for each method for 80k steps with a batch size of 8, using the Adam optimizer (Kingma & Ba, 2017) with the learning rate of 0.0001. During the inference, the length was fixed to 50 tokens for iterative methods.
@@ B.3 Generated Synthetic Sequences
B.4 Additional Synthetic Results
To assess the robustness of BAnG to sequence length, we conducted additional synthetic experiments. Our method was tested on the SingleBind task under three scenarios: sequence lengths of 100 tokens, 200 tokens, and uniformly sampled between 40 and 50 tokens. The results, presented in Table B.2, show that BAnG performs consistently across these variations. The average length of the generated sequences was 46 tokens when trained on sequences sampled from the 40-50 token range.
Table B.2: BAnG performance on SingleBind task with various sequence length. Scores represent the proportion of sequences that contain the correct synthetic motif.
Length Score $\uparrow$
50 0.98
100 0.98
200 0.95
$\mathcal{U}(40,50)$ 0.98
We also investigated BAnG’s behavior when multiple motifs are present simultaneously in a sequence. DualBindMix is a variation of the DualBind task in which two motifs, separated by three tokens, appear in every training sequence, with anchor points placed randomly in the middle of either motif. The task, where motif separation is random, is referred to as DualBindMix Random. The results of these experiments are listed in Table B.3. In the DualBindMix BAnG generated the sequences containing two motifs simultaneously, while in the DualBindMix Random most of the time it generated only one motif. As expected, only the increased uncertainty in motif positioning negatively impacts the performance of our method.
Table B.3: BAnG performance on tasks with several synthetic motifs. Values represent the proportion of sequences containing either one of the synthetic motifs or both.
Task First Second Both
DualBind 0.43 0.53 0.01
DualBindMix 0.01 0.01 0.97
DualBindMix Random 0.47 0.46 0.05
Appendix C Training Data
@@ Appendix D Training Details
As explained in the main text, we conducted the RNA-BAnG training in two steps. In the first step, we used a batch size of 64 and trained for 255k steps. In the second step, we used a batch size of 8 and trained for 216k steps. We stopped training when the validation loss decline became visually negligible. In both training steps, we used the ASMGrad (Reddi et al., 2019) variation of the Adam optimizer (Kingma & Ba, 2017) with default parameters and the learning rate of 0.0001. Learning rate was warmed up linearly for the first 1k steps and then decayed exponentially with $\gamma=0.99$ and a period of 1k steps. Complete training took 4 days on a single MI120 AMD GPU.
Appendix E Evaluation Protocol
@@ E.1 Test Data
The RNA Compendium study provides 244 samples, each comprising a protein sequence paired with approximately 200,000 RNA sequences and their corresponding experimental binding scores. These samples are designated by the authors as RNCMPT00XXX, where XXX represents a numerical identifier. For simplicity, we occasionally refer to these samples solely by their numerical identifiers XXX.
We processed the data as follows. For each sample, RNA sequences were ranked by their binding scores. The top 2,000 sequences were labeled as the positive (interacting) class, while the bottom 2,000 were labeled as the negative (non-interacting) class. To eliminate sequence redundancy and prevent data leakage, duplicate RNA sequences were removed using CD-HIT-EST with a sequence identity threshold of 90%. The remaining sequences were then randomly split into training and testing sets, so that each test set would have a 1,000 sequences. Those test sets are referenced as positive and negative sets in the main manuscript. We then trained individual DeepCLIP models for each sample, following the protocol outlined by the authors. The DeepCLIP model contains approximately 3,000 trainable parameters, and its training protocol – designed by the authors – automatically selects the best-performing weights when provided with sets of positive and negative sequences. Only models achieving an area under the receiver operating characteristic curve (AUROC) of 0.95 or higher on the corresponding test sets were selected for further analysis.
@@ E.2 RNA-BAnG Sampling
For RNA-BAnG, we sample tokens from their distributions using the top-k strategy with $k=4$. Maximum sequence length is set to 50 nucleotides. Average sampling time is 20 minutes for 1,000 RNA sequences. To generate a random nucleotide sequence, we first choose a sequence length at random between 40 and 50 (inclusive). Then, we create a sequence of that length by selecting each nucleotide independently and uniformly at random. The length choice ensures that the mean sequence length of random samples matches that of RNA-BAnG-generated sequences (45.6 nucleotides).
Appendix F Evaluation Results and Comparative Approaches
F.1 Additional Results
@@
We have also compared areas under the threshold-dependent performance curves for GenerRNA and RNAFLOW on corresponding comparison sets (curves shown in the Figure F.1). RNA-BAnG demonstrated area values of 0.81 and 0.53, surpassing GenerRNA with 0.66 and RNAFLOW with 0.13.
Figure F.1: Proportion of sequences above the threshold: The left panel shows results for GenerRNA (yellow), RNA-BAnG (green), and the corresponding positive (blue) and negative (red) experimental sets used in the GenerRNA vs. RNA-BAnG comparison. The right panel presents RNAFLOW (yellow), RNA-BAnG (green), and the positive (blue) and negative (red) experimental sets used in the RNAFLOW vs. RNA-BAnG comparison. The values represent averages over the entire comparison set in each case.
Figure F.2 shows the proportion of high-affinity generated sequences as a function of AlphaFold pLDDT scores and protein sequence similarity with the train set.
@@
Appendix G details protein sequence similarity calculations. Figure F.3 shows kernel density plots calculated over the whole test set of 71 samples. We assessed four different DeepCLIP score threshold values (0.65, 0.75, 0.85, 0.95, from left to right in the Figure). We can conclude that the value of 0.95 is too stringent, as the mean of the positive experimental sets approaches the value of 0.5. Subjectively, the most visually appealing threshold value is 0.75. Nonetheless, in the main text (Figure 6) we continuously assess all the values by plotting the mean proportions above a certain threshold value.
@@
Distributions of DeepCLIP scores for each of 71 test samples are depicted in Figure G.1 and Figure G.2.
F.2 Ablation Studies
To better understand the contributions of individual components in our model, we conducted ablation studies focusing on Geometric Attention and the use of relative position encodings in Cross Attention. Removing the Geometric Attention block resulted in a relatively small change in the number of model parameters (approximately 10%), while removing relative position encodings led to a change of less than 1%.
As shown in Figure F.4, Geometric Attention significantly enhances the performance of RNA-BAnG. While the contribution of relative position encodings in Cross Attention is more modest, it still provides a measurable improvement.
Figure F.4: Proportion of sequences above the threshold: generated by RNA-BAnG (green) and randomly (yellow), the positive (blue) and negative (red) experimental sets. Dashed green line corresponds to results of RNA-BAnG generation without Geometric Attention (on the left) or without relative position encodings in Cross Attention (on the right). The values here represent the averages for the entire comparison set.
F.3 Protein Preprocessing
We also investigated the impact of protein data preprocessing on RNA-BAnG performance. Our findings suggest that segmenting the protein structure into individual domains or removing intrinsically disordered regions (IDRs) can, in some cases, enhance the quality of generated sequences (Figure F.5). By domain here, we mean a compact, spatially distinct region of the protein with an ordered structure — recognized visually, by simply inspecting the structure and seeing how it can be split into distinct rigid blocks.
For instance, in the case of the protein RNCMPT00133 (RNAcompete code), where RNA-BAnG performs the worst, removing IDRs (according to AlphaFold2 pLDDT scores) from the structure led to a tenfold increase in the number of generated high-affinity RNAs. Another example is protein RNCMPT00033, whose AlphaFold2-predicted structure contains multiple domains – running the model on individual domains revealed that some domains yield significantly better results than using the full-length protein. However, such preprocessing is generally only practical when meta-information about protein–RNA interactions is available, as RNAs may interact with multiple domains or IDRs simultaneously.
Figure F.5: Proteins before (on the left) and after (on the right) the preprocessing. The proportion of haRNA generated by RNA-BAnG is reported below the proteins or their domains . The domains were split using visual inspection guided by the pLDDT confidence scores.
F.4 Alternative Sequence Similarity Scoring Approach
We calculated the proportions of highly similar sequences (hsRNA) between the generated and experimental sets. A sequence is considered highly similar if identified as such by blastn with -task blastn parameter. The resulting values for each test sample are shown in Figure F.6. Across different samples, the proportions of generated sequences similar to the positive set are higher than those of similar to the negative set, averaging 29% and 23%, respectively. However, it is important to note the limited applicability of blastn in cases where sequences are intentionally designed without evolutionary signals and are experimentally selected based on the presence of relatively short binding motifs (Ray et al., 2013).
Figure F.6: Proportions of high-similarity RNA sequences, generated by RNA-BAnG, to positive (blue) and negative (red) experimental sets for each test sample. Test samples are ordered by proportion of haRNA sequences, as in Figure 7. Sample IDs mapping to RNAcompete IDs mentioned in Section E.1
F.5 Alternative Motif Search Scoring Approach
Motif searching tools in bioinformatics are specialized software applications designed to identify recurring patterns, or motifs, within biological sequences such as DNA, RNA, or proteins. Motifs are usually represented as Position Weight Matrices (PWMs), which capture the variability and conservation of each nucleotide (or amino acid) at every position within the motif (Stormo et al., 1982). PWMs are obtained by aligning multiple sequences that share the motif and calculating the frequency of each symbol at each position, often followed by converting these frequencies into probabilities or log-odds scores (Schneider & Stephens, 1990). Once constructed, PWMs are then used to scan larger sequences or entire genomes to predict new sites with similar patterns.
As mentioned previously, the sequences in our experimental set are intentionally designed without evolutionary signals, making them inherently random in nature. This randomness makes aligning the sequences problematic. Therefore, we followed the approach used by the RNA Compendium authors, considering only the top-10 scoring 7-mers (scoring data provided by the authors). We then aligned these 7-mers using MAFFT (Katoh & Standley, 2013), removing columns with gaps in more than half of the rows. This removal occurred only at the beginning or end of the alignment, ensuring no changes to the core motifs. The resulting aligned 7-mers were used to construct PWMs for each sample where the data was available in our experimental set.
Figure F.7: Proportion of sequences above the threshold: generated by RNA-BAnG (green) and randomly (yellow), the positive (blue) and negative (red) experimental sets. The values here represent the averages for the entire test set.
Next, we used MOODS (Korhonen et al., 2017) as the motif search tool. We selected it for its ease of use and its ability to incorporate dinucleotide frequencies, which improves search accuracy. We mapped its scores to a 0-1 range using a sigmoid function. Similar to the approach with the DeepCLIP scores, Figure Figure F.7 presents a threshold-dependent performance curve, showing the relationship between sequences with MOODS (sigmoid) scores above a certain threshold and the threshold value itself. A high area value of 0.74 for the positive experimental set, coupled with a very low area value of 0.26 for the negative set, indicates that motif-based scoring is effective at distinguishing high-affinity sequences from others. RNA-BAnG achieved a value of 0.49, indicating moderate success in generating sequences that partially align with the positive set and demonstrating some ability to avoid producing low-affinity outputs. The random set, with an area value of 0.40, suggests that our model outperforms random generation. To provide a more detailed analysis, we defined haRNAs as those in which the highest motif score exceeds 0.4, a threshold selected using the elbow method from Figure F.7. The resulting sample-by-sample statistics are shown in Figure F.8.
This scoring approach correlates with the one based on DeepCLIP scores, showing a Pearson correlation coefficient of 0.64 between the proportion of haRNA sequences defined by DeepCLIP and those defined by motif-based scoring for each sample. Additional details on the usage of MAFFT and MOODS can be found in Appendix G.
Figure F.8: Proportions of high-affinity RNA sequences for each test sample, generated by RNA-BAnG (green) and randomly (yellow). These are compared with the proportions in the positive (blue) and negative (red) experimental sets. Test samples are ordered by RNA-BAnG performance. Sample IDs mapping to RNAcompete IDs mentioned in Section E.1.
Appendix G Alignment Tools Parameters
@@
Clustering of nucleotide sequences was always performed using CD-HIT-EST (Li & Godzik, 2006) with a sequence identity threshold of 90%. We used CD-HIT-EST with the following parameters: -c 0.9 -n 9 -d 0 -T 10 -U 10 -l 9. For protein sequence identity to the train set calculations we used blastp (Altschul et al., 1990) with default parameters. For the nucleotide sequences similarity to the train set calculations we used blastp (Altschul et al., 1990) with default parameters.
For the 7-mers alignment we used MAFFT (Katoh & Standley, 2013) with the --nuc parameter. Motif search was performed with MOODS (Korhonen et al., 2017) using these parameters: -R -p 1.0 --batch. Motifs were represented in adm format. We have also set search window size equal to the motif size.
Figure G.1: Distribution of DeepCLIP scores for the first part of the test samples (RNAcompete sample IDs on top). Scores of RNA-BAnG are in green. Scores of positive and negative experimental sequences are in blue and red, respectively.
Figure G.2: Distribution of DeepCLIP scores for the second part of test samples (RNAcompete sample IDs on top). Scores of RNA-BAnG are in green. Scores of positive and negative experimental sequences are in blue and red, respectively.